OpenStudy (anonymous):

How do i find the number of terms in geometric sequence: 2/81, 4/27, 8/9, ..., 6912?

OpenStudy (anonymous):

a = 2/81 or 0.025 r= 10/81 or .123 tn = ar^n-1 6912 = 0.025 (.123)^n-1 ----- ----- 0.025 0.025 276480 = .123^n-1

OpenStudy (anonymous):

I'm stuck. Many have used the "log" but when do I use that? It doesn't seem to work for this problem..

OpenStudy (zarkon):

8

OpenStudy (zarkon):

\[a_n=\frac{2^n3^{n-1}}{81}\]

OpenStudy (anonymous):

how did you get that? o.O

OpenStudy (zarkon):

\[a_8=6912\]

OpenStudy (anonymous):

a= 2/81 and r = 6 so solve (2/81)(6)^n-1 = 6912

OpenStudy (anonymous):

r = 6? o.o

OpenStudy (anonymous):

no.. it can't be.

OpenStudy (anonymous):

6^n-1 = 6^7 so n=8

OpenStudy (anonymous):

get r by dividing the terms

OpenStudy (zarkon):

\[a_n=\frac{2^n3^{n-1}}{81}=\frac{2\cdot2^{n-1}3^{n-1}}{81}=\frac{2}{81}6^{n-1}\]

OpenStudy (anonymous):

but 2/81 + 6 = 488/81

OpenStudy (anonymous):

it's not in the order of 2/81, 4/27

OpenStudy (anonymous):

ratio you multily don't add

OpenStudy (anonymous):

its geometric!!!

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhh!!!!!

OpenStudy (anonymous):

2/81 x 6 = 4/27

OpenStudy (anonymous):

omg, thank you!!

OpenStudy (anonymous):

I'm so sorry D;

OpenStudy (anonymous):

thats ok